OpenStudy (anonymous):

evaluate limit as x approaches zero from the right. (tan 8x)^x.

OpenStudy (freckles):

\[[\tan(8x)]^x=e^{\ln([\tan(8x)]^x)} \]

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

i would go right to the definition \[\large[\tan(8x)]^x=e^{x\ln([\tan(8x)])}\]

OpenStudy (anonymous):

but what's the answer?

OpenStudy (anonymous):

I think we need L'Hospitals Rule

OpenStudy (freckles):

\[e^{\lim_{x \rightarrow 0^+}x \ln(\tan(8x))}=e^{\lim_{x \rightarrow 0^+}\frac{\ln(\tan(8x))}{\frac{1}{x}}}\]

OpenStudy (anonymous):

The answer was 1

OpenStudy (freckles):

very good

OpenStudy (anonymous):

why didn't you tell me that first??

OpenStudy (freckles):

You have to do the work to determine the answer.